11. ball is thrown vertically upward with a speed of 30.0 m/s.(a) How high does it rise?

 m


(b) How long does it take to reach its highest point?

 s


(c) How long does the ball take to hit the ground after it reaches its highest point?

 s


(d) What is its velocity when it returns to the level from which it started?

 m/s


1
Expert's answer
2022-06-03T15:25:26-0400

Given:

"v_0=30.0\:\rm m/s"

(a)

"h_{\max}=\frac{v_0^2}{2g}=\frac{30.0^2}{2*9.8}=45.9\:\rm m"

(b)

"t=2\frac{v_0}{g}=2*\frac{30.0}{9.8}=6.1\:\rm s"

(c)

"t=\frac{6.1}{2}=3.1\: \rm s"

(d)

"v=v_0=30.0\:\rm s"